Target meeting: 03Feb10, 09:00
Present: JT, CB, PH, JL, KL, EMcC
1. Review of actions:
- PS, JL: write specification for generation of MICE target
signal for ISIS BPS.
- PS: To guide readout and analysis of thermocouple in
Progress has been made, Pierrick has implemented recording. Need to
get data and plot history.
- AD/PH: Will find a way of automating processing so
that BLOC can check the histogram.
- PH: Inform trgt group how archiving of target data
to GRID will be done.
- PS: Look at ways of including the "legal spill" marker
2. Progress on mechanical design and fabrication:
2.1 Target assembly  and : JT
- Pictures circulated using LED lights and tripod for reproducible
position of camera. Overlay of photos does not indicate an increase in
the dust. Possible that the dust arises from the bellows.
- EMcC suggests use of mass spectrometer on the output of the vacuum pump.
But, requires letting up the vacuum. So, we agreed to fit mass
spectrometer in the next iteration of the target test.
- EMcC agreed to find a new window.
- Sunil Patel has tested both Vespel and Torlon. Vespel outgasses
slightly less than the Torlon. One hour and ten hour performance test.
- Alternatively plastics investigated by JT. Plastics with better
CTE and possibly coefficient of friction. Need to make a plan for
materials testing when we've the information on the magnetic analysis.
- Vibrometer system inspected. Advanced instrumentation, but, not
clearly applicable to MICE target because of complexity of MICE target
system. Potentially valuable for QA on vibrations induced by target.
- Parts for two more shafts now ready to go. Lower shaft now has an
integral stop, so is machined from a bar of material. Gun-drilling
company has been contacted. OM has also been contacted re
gun-drilling. Goal is to get two lower shafts with gun-drilling
2.2 Performance analysis of T2#2: PH
- 17 days of operation and despite chiller and UPS failure. In
excess of 1.1M dips and still stable.
- Run for next week to amass 2M dips;
- Then, if the jigs are ready them move on to magnetic measurement,
otherwise keep dipping.
2.3 Stator: CB, GB, OT
- Coil stack and cooling jacket almost complete at Sheffield. Needs
cooling test and then stainless tube will be installed;
- GB has completed design of jig for magnetic measurement.
Goal is to have then manufactured at Imperial over the course of next week
so that they are ready on 22nd February 2010 when Emily returns to RAL and
magnetic measurement could be carried out.
- Stator for T3 will be assembled to point at which weld would have to
be made so that we have T2 and T3 pretty much ready to go in the event
that T1 fails.
Target electronics and DAQ: PS, JL
4. Target paper: CB
- Next phone call: Wednesday; 24Feb10; 09:00
- Date of w/s: Tuesday; 09Mar10, Imperial
- MICE/ISIS target w/s: afternoon of 18Mar10